(8.5x10^10) (3.7x10^5)
1
Expert's answer
2012-06-05T07:50:42-0400
(8.5x10^10)*(3.7x10^5)=8.5*3.7*10^{10+5}=31,45*10^{15}=3.145*10^{16}